OpenStudy (anonymous):

Please Help Solve: 3 ^3square root sign 81 - 2 ^3square root sign 54

OpenStudy (anonymous):

\[3\sqrt[3]{81} - 2\sqrt[3]{54}\]

OpenStudy (anonymous):

find the perfect cube factorisation for each radicand 81=27*3 54=27*2

hero (hero):

54 is not a cube, is it?

hero (hero):

Oh nevermind

OpenStudy (anonymous):

thanks!

hero (hero):

You still won't be able to subtract anything though

OpenStudy (anonymous):

this one too:

OpenStudy (anonymous):

\[=9\sqrt[3]{3}-6\sqrt[3]{2}\]

OpenStudy (anonymous):

\[\sqrt[4]{32} + \sqrt[4]{48}\]

OpenStudy (anonymous):

32=16*2; 48=16*3 and 16 is 2^4

OpenStudy (anonymous):

\[=\sqrt[4]{16*2}+\sqrt[4]{16*3}=2\sqrt[4]{2}+2\sqrt[4]{3}\]

OpenStudy (anonymous):

thanks for the help!

OpenStudy (anonymous):

welcome!